# From Parameters to Behaviors: Unsupervised Compression of the Policy Space

### v1

@@ 1 Introduction
High-dimensional parameterization of policies via deep neural networks has been a key driver of recent successes in Deep Reinforcement Learning (among others, Andrychowicz et al., 2020; Smith et al., 2022; Bakhtin et al., 2022; Wurman et al., 2022; Duval et al., 2024). A major drawback of this approach, however, is a significant increase in sample complexity, which is further compounded when the agent is called to solve multiple and potentially unknown tasks, typically requiring learning tabula rasa (Agarwal et al., 2022). This inefficiency often stems from a fundamental redundancy in the parameter space, where a large set of distinct weight configurations maps to a much smaller set of effective behaviors. Various approaches tried to solve this limitation as a byproduct, such as explicitly learning diverse behaviors (Eysenbach et al., 2018; Zahavy et al., 2022; De Paola et al., 2025; Zamboni et al., 2025), or enforcing small policy networks in asymmetric actor-critic architectures (Duval et al., 2024; Mastikhina et al., 2025).
In this paper, we address this limitation directly, through the lenses of the Manifold Hypothesis (Cayton et al., 2005), a widely accepted tenet of Machine Learning, and we hypothesize that it holds in RL as well, namely that:
@@
In view of this hypothesis, we propose a paradigm shift from learning in the parameter space to learning in the (latent) behavior space itself. To do so, the agent first needs to learn a latent representation of the possible behaviors, which, according to the aforementioned hypothesis, should be low-dimensional and policy network invariant. Then, it needs to find a way to leverage this representation to solve different tasks efficiently. The proposed solution is a novel two-stage framework directly inspired by the Unsupervised RL formalism (Laskin et al., 2021), allowing for the explicit exploitation of this latent structure. In a first pre-training phase, we learn a latent representation of the behavior manifold by leveraging a generative model in a fully unsupervised fashion, that is, without including any information related to a specific task, i.e., reward. In this way, we can learn a latent structure that models the intrinsic nature of the environment dynamics, rather than its coupling with a task, and preserve the end-to-end differentiability that makes gradient-based optimization effective. In a second fine-tuning phase, we leverage the pre-trained representation to fine-tune policies against specific tasks known a posteriori, avoiding the need to learn from scratch. In particular, the fine-tuning phase involves performing gradient steps in the latent space, thereby optimizing latent behaviors directly. This approach enables the agent to explore the inherently low-dimensional behavior space rather than the high-dimensional parameter space.
@@
Research Questions: (Q1) Is it possible to learn a low-dimensional latent representation of high-dimensional policy parameter spaces in an unsupervised fashion? (Q2) What are the properties of such a latent representation, if it exists? Is its intrinsic dimension a function of the behavioral complexity rather than the size of the parameter space? (Q3) How can we fine-tune against specific tasks leveraging the low-dimensional space? Does this come with positives?
Content Outline and Contributions.  First, in Section 3, we formulate the problem of learning a latent representation of behaviors in an unsupervised fashion and then leveraging it to solve specific tasks. Then, in Section 4, we characterize our proposed solution to such a problem, namely, to address it in a two-stage pipeline. Finally, in Section 5, we perform experiments extensively to address the Research Questions. We show that, indeed, the proposed pipeline is able to learn low-dimensional latent representations (Q1), which are more influenced by the environment than by the size of the compressed policies (Q2), and, finally, that learning over this reduced space can make simple algorithms competitive against complex state-of-the-art DRL algorithms (Q3).
@@ 2 Preliminaries
Notation.  In the following, we denote a set with a calligraphic letter $\mathcal{A}$} and its size as $|{\mathcal{A}}|$, the simplex on $\mathcal{A}$} is denoted as $\Delta({\mathcal{A}}):=\{p\in[0,1]^{|{\mathcal{A}}|}\mid\sum_{a\in{\mathcal{A}}}p(a)=1\}$. For two distributions $p_{1},p_{2}\in\Delta({\mathcal{A}})$, we define a general measure of divergence between distributions with $D(p_{1}||p_{2})$.
@@
Interaction Protocol.  As a base model for interaction, we consider a (finite-horizon) Controlled Markov Process (CMP). A CMP is defined as the tuple ${\mathcal{M}}:=({\mathcal{S}},{\mathcal{A}},{\mathbb{P}},\mu,T)$, where $\mathcal{S}$} is the state space and $\mathcal{A}$} is the action space. At the start of an episode, the initial state $s_{0}$ of $\mathcal{M}$} is drawn from an initial state distribution $\mu\in\Delta({\mathcal{S}})$. Upon observing $s_{0}$, the agent takes action $a_{0}\in{\mathcal{A}}$, the system transitions to $s_{1}\sim{\mathbb{P}}(\cdot\mid s_{0},a_{0})$ according to the transition model ${\mathbb{P}}:{\mathcal{S}}\times{\mathcal{A}}\to\Delta({\mathcal{S}})$. The process is repeated until $T$ is reached and $s_{T}$ is generated, being $T<\infty$ the horizon of an episode. The agent selects actions according to a decision policy $\pi:{\mathcal{S}}\to\Delta({\mathcal{A}})$ such that $\pi(a|s)$ denotes the conditional probability of taking action $a$ upon observing state $s$. Deploying a policy $\pi$ over $\mathcal{M}$} leads to the generation of trajectories $\tau$, defined as a sequence of state-action pairs $\tau:=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{T})$. Furthermore, a policy $\pi$ induces a state distribution $d_{\pi}^{s}\in\Delta({\mathcal{S}})$ over the state space of the CMP $\mathcal{M}$} defined as $d_{\pi}^{s}(s)=\sum_{t=0}^{T}Pr(s_{t}=s)$. It also induces a state-action distribution $d_{\pi}^{sa}\in\Delta({\mathcal{S}}\times{\mathcal{A}})$, defined as $d_{\pi}^{sa}(s,a)=\pi(a\mid s)d_{\pi}^{s}(s)$, which we will denote as behaviors. In the following, we will consider deterministic policies $\pi_{\bm{\theta}}:{\mathcal{S}}\to{\mathcal{A}}$ represented by neural networks parametrized by a set of weights ${\bm{\theta}}\in\Theta$, where $\Theta\subseteq{\mathbb{R}}^{P}$ is the policy parameter space, with $P$ being the total number of parameters. We will denote with $\Pi_{\Theta}$ the Policy Space, namely the collection of the policies that such parameters can represent. For brevity of notation, we define a policy $\pi_{\bm{\theta}}$ as its set of parameters $\bm{\theta}$} and the policy space $\Pi_{\Theta}$ as the parameter space $\Theta$ that induces it.
@@
(Unsupervised) RL.  In RL, an agent learns how to solve (downstream) tasks, encoded by different reward signals. For this matter, we define a Markov Decision Process (MDP, Puterman, 2014) ${\mathcal{M}}_{R}:={\mathcal{M}}\cup R$ as a coupling of a CMP $\mathcal{M}$} and a reward function $R:{\mathcal{S}}\times{\mathcal{A}}\to{\mathbb{R}}$, which the agent observes after every state transition. In the Unsupervised Reinforcement Learning (URL, Laskin et al., 2021) framework, the reward signal is not always available to the agent from the beginning. It often belongs to a (potentially infinite) family of tasks $\mathcal{R}$}, also unknown to the agent. URL is then composed of two phases: (1) an unsupervised pre-training phase involves the agent interacting with a CMP to acquire general-purpose knowledge without receiving any reward signal, which is distilled into a pre-trained model $\mathbb{M}$}; (2) the supervised fine-tuning phase begins once a reward function $R\in{\mathcal{R}}$ is revealed. At this point, the CMP becomes a standard MDP ${\mathcal{M}}_{R}$, and the agent leverages the pre-trained model $\mathbb{M}$} to find a set of policy parameters that maximizes the expected return for the given task, namely as
${\bm{\theta}}^{*}=\operatorname*{arg\,max}_{{\bm{\theta}}\in\Theta}J^{R}({\bm{\theta}},{\mathbb{M}})=\operatorname*{arg\,max}_{{\bm{\theta}}\in\Theta}\mathbb{E}_{(s,a)\sim d_{\pi_{\bm{\theta}}}^{sa},{\mathbb{M}}}[R(s,a)].$ (1)
@@
Policy Optimization.  Directly optimizing against the set of policy parameters, a process called Policy Optimization (PO, Deisenroth et al., 2013), gives quite surprising results. In particular, when deep neural policies are considered, first-order methods have been extensively employed, with a popular approach to PO being Policy Gradient (PG, Peters & Schaal, 2008), which updates the parameters by simple gradient ascent ${\bm{\theta}}^{\prime}\leftarrow{\bm{\theta}}+\alpha\nabla_{\bm{\theta}}J({\bm{\theta}})$. Among others, Policy Gradient with Parameter-based Exploration (PGPE, Sehnke et al., 2008) is a PG algorithm that handles exploration in the parameter space by sampling the policy parameters $\bm{\theta}$} from a hyper-policy $\nu_{\bm{\phi}}$, parametrized by $\bm{\phi}$}. [^1] For instance, Gaussian hyper-policies will be parametrized by their mean and standard deviation, i.e. ${\bm{\phi}}=(\mu,\sigma).$ PGPE optimizes a trajectory-based version of the objective defined in Eq. 1, defined as:
$J^{R}({\bm{\theta}},{\bm{\phi}},{\mathbb{M}})=\mathbb{E}_{\tau\sim p(\cdot\mid{\bm{\theta}}),{\bm{\theta}}\sim\nu_{\bm{\phi}},{\mathbb{M}}}[R(\tau)],$ (2)
@@
where $R(\tau)=\sum_{t=0}^{T}R(s_{t},a_{t})$ is the return of a trajectory, and $p(\tau\mid{\bm{\theta}})=\mu(s_{0})\prod_{t=0}^{T}{\mathbb{P}}(s_{t+1}\mid s_{t},a_{t})\pi_{\bm{\theta}}(a_{t}|s_{t})$ is the probability density of a trajectory. In PGPE, the parameter vector $\bm{\phi}$} is usually updated via gradient ascent using a Monte Carlo estimator of the gradient computed over $N\in{\mathbb{N}}$ trajectories:
@@
Generative Models.  Generative models have achieved remarkable success in density estimation for multi-modal data, drawing significant interest from the RL community. Among others, Autoencoders (AE, Hinton & Salakhutdinov, 2006) are a type of artificial neural network used to learn efficient data encoding in an unsupervised manner. The aim is first to learn encoded representations of data and then generate the input data (as closely as possible) from the learned encoded representations. More specifically, their goal is to map a data space ${\mathcal{X}}\subseteq{\mathbb{R}}^{n}$ to a latent space ${\mathcal{Z}}\subseteq{\mathbb{R}}^{k}$, with $k\ll n$. AEs are composed of an encoder, a function $f_{\bm{\xi}}:{\mathcal{X}}\to{\mathcal{Z}}$, parametrized by vector $\bm{\xi}$}, which maps a data sample $x\in{\mathcal{X}}$ to a latent code $z\in{\mathcal{Z}}$, and a decoder, a function $g_{\bm{\zeta}}:{\mathcal{Z}}\to{\mathcal{X}}$, parametrized by vector $\bm{\zeta}$}, which reconstructs the data sample $\hat{x}\in{\mathcal{X}}$ from the latent code $z$ in such a way that $g_{\bm{\zeta}}\approx f_{\bm{\xi}}^{-1}$. An AE is typically trained by minimizing the reconstruction error ${\mathcal{L}}_{AE}({\bm{x}})=d({\bm{x}},g_{\bm{\zeta}}(f_{\bm{\xi}}({\bm{x}})))$, where $d$ is a metric that measures the distance of the samples in the data space. These sorts of architectures are particularly compelling in view of the Manifold Hypothesis (Cayton et al., 2005): AEs learn this underlying structure by compressing the data into a compact latent space that represents the manifold and then reconstructing the original data from it, as illustrated in Fig. 1. Unfortunately, AEs are far from being bulletproof.
@@
In cases where no plausible embedding exists, even networks $(f_{\bm{\xi}},g_{\bm{\zeta}})$ which come close to perfectly reconstructing the manifold $\mathcal{M}$ will incur numerical instability (Cornish et al., 2020). In some other cases, it is possible to resolve these topological issues by increasing the latent dimension $k$. For instance, a dimensionality of $k=2d^{\star}+1$ is enough to topologically embed any manifold of dimension $d^{\star}$ in $\mathbb{R}$ (Theorem V3, Hurewicz & Wallman, 2015).
@@ 3 Problem Formulation
$\forall{\bm{\theta}}\in\Theta,\quad\exists{\bm{z}}\in{\mathcal{Z}}:\quad g^{\star}=\operatorname*{arg\,min}_{g}D(d_{\pi_{\bm{\theta}}}^{sa}||d_{\pi_{g({\bm{z}})}}^{sa}).$ (4)
This task is essentially unsupervised, as any notion of a specific task is absent. Indeed, it is somewhat similar to the Policy Space Compression framework (Mutti et al., 2022), yet in the latter, the authors seek a way to reduce the cardinality of the policy space, rather than its dimensionality. Moreover, the constraints defining a valid compression are stricter than ours, resulting in an optimization problem that is known to be NP-hard.
Once such a low-dimensional space and generative function are available, solving for different tasks will call for simply searching over a simpler space than the original one, which we call Latent Behavior Optimization. In other words, the standard PO problem of Eq. 1, namely of finding an optimal policy parameter vector ${\bm{\theta}}^{*}\in\Theta$ will be reformulated as the problem of finding an optimal latent code ${\bm{z}}^{*}\in{\mathcal{Z}}$. For a given task with reward $R\in{\mathcal{R}}$, the policy optimization problem is now defined as:
${\bm{z}}^{*}=\operatorname*{arg\,max}_{{\bm{z}}\in{\mathcal{Z}}}J^{R}({\bm{z}})=\operatorname*{arg\,max}_{{\bm{z}}\in{\mathcal{Z}}}J^{R}({\bm{\theta}}=g({\bm{z}})).$ (5)
On the other hand, this task is essentially supervised, as it is well-defined as soon as the agent is provided with a reward. In the following, we will show how the URL framework can indeed provide essential tools in addressing the two problems.

### v2

@@ 1 Introduction
High-dimensional parameterization of policies via deep neural networks has been a key driver of recent successes in Deep Reinforcement Learning (among others, Andrychowicz et al., 2020; Smith et al., 2022; Bakhtin et al., 2022; Wurman et al., 2022; Degrave et al., 2022). A major drawback of this approach, however, is a significant increase in sample complexity, which is further compounded when the agent is called to solve multiple and potentially unknown tasks, typically requiring learning tabula rasa (Agarwal et al., 2022). This inefficiency often stems from a fundamental redundancy in the parameter space, where a large set of distinct weight configurations maps to a much smaller set of effective behaviors. Various approaches tried to solve this limitation as a byproduct, such as explicitly learning diverse behaviors (Eysenbach et al., 2018; Zahavy et al., 2023; De Paola et al., 2025; Zamboni et al., 2025), or enforcing small policy networks in asymmetric actor-critic architectures (Degrave et al., 2022; Mastikhina et al., 2025).
In this paper, we address this limitation directly through the lens of the Manifold Hypothesis (Cayton et al., 2005), a widely accepted tenet of Machine Learning, and we hypothesize that it holds in Reinforcement Learning as well, namely that:
@@
In view of this hypothesis, we propose a paradigm shift from learning in the parameter space to learning in the (latent) behavior space itself. To do so, the agent first needs to learn a latent representation of the possible behaviors, which, according to the aforementioned hypothesis, should be low-dimensional and policy network invariant. Then, it needs to find a way to leverage this representation to solve different tasks efficiently. The proposed solution is a novel two-stage framework directly inspired by the Unsupervised RL formalism (Laskin et al., 2021), allowing for the explicit exploitation of this latent structure. In a first pre-training phase, we learn a latent representation of the behavior manifold by leveraging a generative model in a fully unsupervised fashion, that is, without including any information related to a specific task, i.e., reward. In this way, we can learn a latent structure that models the intrinsic nature of the environment dynamics, rather than its coupling with a task, and preserve the end-to-end differentiability that makes gradient-based optimization effective. In a second fine-tuning phase, we leverage the pre-trained representation to fine-tune policies on specific tasks known a posteriori, avoiding the need to learn from scratch. In particular, the fine-tuning phase involves performing gradient steps in the latent space, thereby optimizing latent behaviors directly. This approach enables the agent to explore the inherently low-dimensional behavior space rather than the high-dimensional parameter space.
@@
Research Questions: (Q1) Is it possible to learn a low-dimensional latent representation of a high-dimensional policy parameter space in an unsupervised fashion? (Q2) What are the properties of such a latent representation, if it exists? Is its intrinsic dimension a function of the behavioral complexity rather than the size of the parameter space? (Q3) How can we fine-tune on specific tasks leveraging the low-dimensional space? Does this offer any advantages?
Content Outline and Contributions.  First, in Section 3, we formulate the problem of learning a latent representation of behaviors in an unsupervised fashion and then leveraging it to solve specific tasks. Then, in Section 4, we characterize our proposed solution to this problem, namely, addressing it in a two-stage pipeline. Finally, in Section 5, we perform experiments extensively to address the Research Questions. We demonstrate that the proposed pipeline is indeed able to learn low-dimensional latent representations (Q1), which are more influenced by the environment than by the size of the compressed policies (Q2). Finally, we demonstrate that learning over this reduced space can make simple algorithms competitive against complex state-of-the-art DRL algorithms (Q3).
@@ 2 Preliminaries
Notation.  In the following, we denote a set with a calligraphic letter $\mathcal{A}$} and its size as $|{\mathcal{A}}|$. For a finite set, $\Delta({\mathcal{A}})$ denotes the probability simplex. For continuous spaces, we denote the set of Borel probability measures as $\mathcal{P}({\mathcal{A}})$. When the context allows, we use $\Delta({\mathcal{A}})$ to refer generally to the space of valid distributions over $\mathcal{A}$}. For two distributions $p_{1},p_{2}\in\Delta({\mathcal{A}})$, we define a general measure of divergence between distributions with $D(p_{1}\parallel p_{2})$.
@@
Interaction Protocol.  As a base model for interaction, we consider a (finite-horizon) Controlled Markov Process (CMP). A CMP is defined as the tuple ${\mathcal{M}}\coloneq({\mathcal{S}},{\mathcal{A}},{\mathbb{P}},\mu,T)$, where $\mathcal{S}$} is the state space and $\mathcal{A}$} is the action space. At the start of an episode, the initial state $s_{0}$ of $\mathcal{M}$} is drawn from an initial state distribution $\mu\in\Delta({\mathcal{S}})$. Upon observing $s_{0}$, the agent takes action $a_{0}\in{\mathcal{A}}$, and the system transitions to $s_{1}\sim{\mathbb{P}}(\cdot\mid s_{0},a_{0})$ according to the transition model ${\mathbb{P}}:{\mathcal{S}}\times{\mathcal{A}}\to\Delta({\mathcal{S}})$. The process is repeated until $T$ is reached and $s_{T}$ is generated, with $T<\infty$ being the horizon of an episode. The agent selects actions according to a decision policy $\pi:{\mathcal{S}}\to\Delta({\mathcal{A}})$ such that $\pi(a\mid s)$ denotes the conditional probability of taking action $a$ upon observing state $s$. Deploying a policy $\pi$ over $\mathcal{M}$} leads to the generation of trajectories $\tau$, defined as a sequence of state-action pairs $\tau:=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{T})$. Furthermore, a policy $\pi$ induces a state distribution $d_{\pi}^{s}\in\Delta({\mathcal{S}})$ over the state space of the CMP $\mathcal{M}$} defined as $d_{\pi}^{s}(s)=\frac{1}{T+1}\sum_{t=0}^{T}\Pr(s_{t}=s)$. It also induces a state-action distribution $d_{\pi}^{sa}\in\Delta({\mathcal{S}}\times{\mathcal{A}})$, defined as $d_{\pi}^{sa}(s,a)=\pi(a\mid s)\ d_{\pi}^{s}(s)$, which we will denote as the behavior of the policy. In the following, we will consider deterministic policies $\pi_{\bm{\theta}}:{\mathcal{S}}\to{\mathcal{A}}$ represented by neural networks parameterized by a set of weights ${\bm{\theta}}\in\Theta$, where $\Theta\subseteq{\mathbb{R}}^{P}$ is the policy parameter space, with $P$ being the total number of parameters. We define the Policy Space $\Pi_{\Theta}$ as the collection of policies that can be represented by $\Theta$. For brevity of notation, we denote a policy $\pi_{\bm{\theta}}$ as its set of parameters $\bm{\theta}$} and the policy space $\Pi_{\Theta}$ as the parameter space $\Theta$ that induces it.
@@
(Unsupervised) RL.  In RL, an agent learns how to solve (downstream) tasks, encoded by different reward signals.To this end, we define a Markov Decision Process (MDP, Puterman, 2014) ${\mathcal{M}}_{R}:={\mathcal{M}}\cup R$ as a coupling of a CMP $\mathcal{M}$} and a reward function $R:{\mathcal{S}}\times{\mathcal{A}}\to{\mathbb{R}}$, which the agent observes after every state transition. In the Unsupervised Reinforcement Learning (URL, Laskin et al., 2021) framework, the reward signal is not always available to the agent from the beginning. It often belongs to a (potentially infinite) family of tasks $\mathcal{R}$}, also unknown to the agent. URL is then composed of two phases: (1) an unsupervised pre-training phase involves the agent interacting with a CMP to acquire general-purpose knowledge without receiving any reward signal, which is distilled into a pre-trained model $\mathbb{M}$}; (2) the supervised fine-tuning phase begins once a reward function $R\in{\mathcal{R}}$ is revealed. At this point, the CMP becomes a standard MDP ${\mathcal{M}}_{R}$, and the agent leverages the pre-trained model $\mathbb{M}$} to find a set of policy parameters that maximizes the expected return for the given task, namely as:
${\bm{\theta}}^{*}=\operatorname*{arg\,max}_{{\bm{\theta}}\in\Theta}J^{R}({\bm{\theta}},{\mathbb{M}})=\operatorname*{arg\,max}_{{\bm{\theta}}\in\Theta}\mathbb{E}_{(s,a)\sim d_{\pi_{\bm{\theta}}}^{sa},{\mathbb{M}}}\left[R\left(s,a\right)\right].$ (1)
@@
Policy Optimization.  Policy Optimization (PO, Deisenroth et al., 2013), which involves optimizing the policy parameters directly, has shown surprisingly good results. This is especially true for deep neural policies, where first-order methods have been extensively employed. A popular approach to PO is Policy Gradient (PG, Peters & Schaal, 2008), which updates the parameters by simple gradient ascent ${\bm{\theta}}^{\prime}\leftarrow{\bm{\theta}}+\alpha\nabla_{\bm{\theta}}J({\bm{\theta}})$. Among others, Policy Gradient with Parameter-based Exploration (PGPE, Sehnke et al., 2008) is a PG algorithm that handles exploration in the parameter space by sampling the policy parameters $\bm{\theta}$} from a hyper-policy $\nu_{\bm{\phi}}$, parameterized by $\bm{\phi}$}. [^1] For instance, Gaussian hyper-policies will be parameterized by their mean and standard deviation, i.e. ${\bm{\phi}}=({\bm{\mu}},{\bm{\sigma}}).$ PGPE optimizes a trajectory-based version of the objective defined in Eq. 1, defined as:
$J^{R}({\bm{\theta}},{\bm{\phi}},{\mathbb{M}})=\mathbb{E}_{\tau\sim p\left(\cdot\;\middle|\;{\bm{\theta}}\right),{\bm{\theta}}\sim\nu_{\bm{\phi}},{\mathbb{M}}}[R(\tau)],$ (2)
@@
where $R(\tau)=\sum_{t=0}^{T}R(s_{t},a_{t})$ is the return of a trajectory, and $p(\tau\mid{\bm{\theta}})=\mu(s_{0})\prod_{t=0}^{T}{\mathbb{P}}(s_{t+1}\mid s_{t},a_{t})\ \pi_{\bm{\theta}}(a_{t}\mid s_{t})$ is the probability density of a trajectory. In PGPE, the parameter vector $\bm{\phi}$} is usually updated via gradient ascent using a Monte Carlo estimator of the gradient computed over $N\in{\mathbb{N}}$ trajectories:
@@
Generative Models.  Generative models have achieved remarkable success in density estimation for multi-modal data, drawing significant interest from the RL community. Among others, Autoencoders (AE, Hinton & Salakhutdinov, 2006) are a type of artificial neural network used to learn efficient data encoding in an unsupervised manner. The aim is first to learn encoded representations of data and then generate the input data (as closely as possible) from the learned encoded representations. More specifically, their goal is to map a data space ${\mathcal{X}}\subseteq{\mathbb{R}}^{n}$ to a latent space ${\mathcal{Z}}\subseteq{\mathbb{R}}^{k}$, with $k\ll n$. AEs are composed of an encoder, a function $f_{\bm{\xi}}:{\mathcal{X}}\to{\mathcal{Z}}$, parameterized by vector $\bm{\xi}$}, which maps a data sample ${\bm{x}}\in{\mathcal{X}}$ to a latent code ${\bm{z}}\in{\mathcal{Z}}$, and a decoder, a function $g_{\bm{\zeta}}:{\mathcal{Z}}\to{\mathcal{X}}$, parameterized by vector $\bm{\zeta}$}, which reconstructs the data sample $\hat{\bm{x}}\in{\mathcal{X}}$ from the latent code $\bm{z}$} in such a way that $g_{\bm{\zeta}}\approx f_{\bm{\xi}}^{-1}$. An AE is typically trained by minimizing the reconstruction error ${\mathcal{L}}_{AE}({\bm{x}})=d({\bm{x}},(g_{\bm{\zeta}}\circ f_{\bm{\xi}})({\bm{x}}))$, where $d$ is a metric that measures the distance of samples in the data space. These architectures are particularly compelling in light of the Manifold Hypothesis (Cayton et al., 2005): AEs learn the underlying structure by compressing the data into a compact latent space that represents the manifold and then reconstructing the original data from it, as illustrated in Fig. 1.
@@
Unfortunately, AEs are far from being bulletproof. In cases where no plausible embedding exists, even networks $(f_{\bm{\xi}},g_{\bm{\zeta}})$ which come close to perfectly reconstructing the manifold $\mathcal{M}$ will incur numerical instability (Cornish et al., 2020). In some other cases, it is possible to resolve these topological issues by increasing the latent dimension $k$. For instance, a dimensionality of $k=2d^{\star}+1$ is enough to topologically embed any manifold of dimension $d^{\star}$ in $\mathbb{R}^{k}$ (Theorem V3, Hurewicz & Wallman, 2015).
@@ 3 Problem Formulation
$\forall{\bm{\theta}}\in\Theta,\quad\exists{\bm{z}}\in{\mathcal{Z}}:\quad g^{\star}=\operatorname*{arg\,min}_{g}D\left(d_{\pi_{\bm{\theta}}}^{sa}\parallel d_{\pi_{{\bm{z}}}}^{sa}\right).$ (4)
This task is essentially unsupervised, as any notion of a specific task is absent. Indeed, it is somewhat similar to the Policy Space Compression framework (Mutti et al., 2022), yet in the latter, the authors aim to reduce the cardinality of the policy space, rather than its dimensionality. Moreover, the constraints defining a valid compression are stricter than ours, resulting in an optimization problem that is known to be NP-hard.
Once such a low-dimensional space and generative function are available, solving for different tasks will require searching over a simpler space than the original one. We call this process Latent Behavior Optimization. In other words, the standard PO problem of Eq. 1, which requires finding an optimal policy parameter vector ${\bm{\theta}}^{*}\in\Theta$, will be reformulated as the problem of finding an optimal latent code ${\bm{z}}^{*}\in{\mathcal{Z}}$ that, via the generative function $g$, yields $\bm{\theta}$$^{*}$. For a given task with reward $R\in{\mathcal{R}}$, the policy optimization problem is now defined as:
${\bm{z}}^{*}=\operatorname*{arg\,max}_{{\bm{z}}\in{\mathcal{Z}}}J^{R}({\bm{z}})=\operatorname*{arg\,max}_{{\bm{z}}\in{\mathcal{Z}}}J^{R}\left({\bm{\theta}}=g\left({\bm{z}}\right)\right).$
Contrary to the Latent Behavior Compression task, this task is essentially supervised, as it is well-defined as soon as the agent is provided with a reward. In the following, we will show how the URL framework can indeed provide essential tools in addressing the two problems.